NEW YORK JOURNAL OF MATHEMATICS CONFERENCE JUNE 8–15, 1997

LIST OF PARTICIPANTS

Ademi Adewuyi

Department of Mathematics, Vrije Universitiet, Brussel, Belgium mw951257@beta.ufsia.ac.be

John Albert

Department of Mathematics, Univ of Oklahoma, Norman, OK 73019 jalbert@ou.edu

Idris Assani

Department of Mathematics, Univ of Norht Carolina, Chapel Hill, NC 27599 assani@math.unc.edu

Dan Barbasch

Department of Mathematics, White Hall, Cornell Univ, Ithaca, NY 14853 barbasch@cornell.edu

Olga Beaver

Department of Mathematics, Williams College, Williamstown, MA 01267 Olga.R.Beaver@williams.edu

A. Berenstein

Department of Mathematics, Cornell University

J. Bona

Texas Institute for Computational and Applied Mathematics, The University of Texas at Austin, Austin, TX 78712 bona@math.utexas.edu

Rob Bradley:

Department of Mathematics, Adelphi Univ,Garden City, NY 11530 Bradley@panter.adelphi.edu Luis Casian Department of Mathematics, Ohio State Univ., Columbus, OH 43210 casian@mathserv.mps.ohio-state.edu

Natalya Chekhova Natalya Chekhova, etudiants - equipe SDD, IML, Case 930, 163 av. de Luminy, F13288 Marseille Cedex 9, FRANCE nata@iml.univ-mrs.fr

Hongqiu Chen Department of Mathematics, University of Texas at Austin, Austin, TX 78712

J-P Jeff Chen Department of Mathematics, Yale Univ., New Heaven, CT 06520 jpchen@math.yale.edu

Jal Choksi Department of Mathematics, McGill University, Montreal, Quebec, Canada Choksi@math.mcgill.ca

Walter Craig Mathematics Department, Brown University, Providence, RI 02912 craigw@math.brown.edu

Burgess Davis Department of Mathematics, Purdue University, bdavis@stat.purdue.edu

Karma Dajani dajani@math.ruu.nl

Andres del Junco Department of Mathematics, University of Toronto, Toronto, Ontario, Canada deljunco@math.toronto.edu

Sakin Demir Department of Mathematics, Univ of Illinois, Urbana, IL 61801 s-demir1@math.uiuc.edu Stanley Eigen Dep. of Mathematics, Noreastern Univ., Boston, MA eigen@neu.edu

Sam Evens Department of Mathematics, Univ. of Arizona, Tucson, AZ 85721 evens@math.arizona.edu

Sebastien Ferenczi IML, Case 930, 163 av. de Luminy, F13288 Marseille Cedex 9, FRANCE ferenczi@iml.univ-mrs.fr

Sebastian Ferrando Department of Mathematics, Ryerson University Toronto, Canada Ferrando@acs.ryerson.ca

Adam Fieldsteel Department of Mathematics, Wesleyan University, Middletown, CT 06459 afieldsteel@math.wesleyan.edu

Nat Friedman Department of Mathematics, SUNY-University at Albany, Albany, NY 12222 artmath@math.albany.edu

Maria Fung Department of Mathematics, Cornell Univ, Ithaca, NY 14853 gargova@math.cornell.edu

W. Graham Institute of Advanced Studies, Princeton

Daniel Goldstein, Univ of Vermont 199 Sandra Circle Burlinghton, Vermont 05401 dgoldste@math.uvm.edu

Nader Goubran Department of Mathematics, Univ of Illinois, Urbana, IL Toshihiro Hamachi Department of Mathematics, Kyushhu University, Japan d75439a@kyu-cc.cc.kyushu-u.ac.jp

Jane Hawkins Department of Mathematics, Univ of Norht Carolina, Chapel Hill, NC 27599 jmh@math.unc.edu

Martin Hildebrand Department of Mathematics, SUNY-University at Albany, Albany, NY 12222 martinhi@math.albany.edu

Marc Jaffrey Department of Mathematics, University at Albany, Albany, NY 12222 mj3109@csc.albany.edu

Udo Hertrich-Jeromin Department of Mathematics, Univ of Massachusetts, Amherst, MA 01003 jeromin@math.umass.edu

R. Howe Department of Mathematics, Yale University

Chaocheng Huang Department of Mathematics, Wright State Univ, Dayton, OH 45435 chuang@gauss.math.wright.edu

Lizhen Ji Department of Mathematics, University of Michigan, Ann Arbor

Dihua Jiang Department of Mathematics, Yale Univ, New Heaven, CT 06520 dhjiang@math.yale.edu

Aimee Johnson Department of Mathematics, Swarthmore College, Swarthmore, PA 19081 aimee@swarthmore.edu Roger L. Jones Department of Mathematics, DePaul University, Chicago, IL rjones@condor.depaul.edu

Julee Kim, Yale Univ. P.O. Box 208283, Department of Mathematics, Yale Station, New Heaven, CT 06520 julee@pascal.math.yale.edu

Isaac Kornfeld Department of Mathematics, North Dakota State Univ, Fargo, ND 58105 kornfeld@plains.nodak.edu

B. Kostant Department of Mathematics, M.I.T.

Kunquan Lan Department of Mathematics, Univ. of Glasgow, Glasgow, G12 8QW, Scotland, UK kl@maths.gla.ac.uk

Jian–Shu Li Department of Mathematics, University of Maryland,

Yue Liu Brown University and University of Texas at Arlington

Garry Morris Department of Mathematics, Univ. East Anglia, Norwich, NR6 5LB, England G.Morris@uea.ac.uk

Radhakrishnan Nair Mathematical Sciences, University of Liverpool, Liverpool L69 3BX, United Kingdom nair@liv.ac.uk

Kenneth Nwabueze MSRI, 11500 Menlo Ave, Hawthorne, CA nwabueze@msri.org Jim Olsen Department of Mathematics, North Dakota State Univ, Fargo, ND 58105 jolsen@plains.nodak.edu

Jan Paradowski Department of Mathematics, LGRT Box 34515, University of Massachusetts at Amherst, Amherst, MA 01003 jp@lie.math.umass.edu

Kyewon Park Department of Mathematics, Ajou University, Suwon 442-749, Korea kkpark@madang.ajou.ac.kr

Raj Prasad Department of Mathematics, University of Lowell,, MA, PRASADV@woods.uml.edu

A. Postnikov Department of Mathematics, M.I.T.

Mahbubur Rahman Department of Mathematics, Concordia Univ. Canada rahman1@vax2.concordia.ca

Karin Reinhold Department of Mathematics, University at Albany, SUNY, Albany, NY 12222 reinhold@math.albany.edu

Joe Rosenblatt Department of Mathematics, Illinois University, Urbana, IL jrsnbltt@math.uiuc.edu

Juergen Rohlfs Katholische Universitaet Eichstaett, Ostenstr. 26, 85072 Eichstaett, Germany juergen.rohlfs@ku-eichstaett.de S. Sahi Department of Mathematics, Rutgers University

Ayse Sahin Department of Mathematics, North Dakota State Univ, Fargo, ND 58105 sahin@plains.nodak.edu

Shayan Sen Department of Mathematics, Cornell Univ, Ithaca, NY 14853 shayan@math.cornell.edu

Mark Sepanski Department of Mathematics, Oklahoma State Univ, 401 Mathematics Science, Stillwater, OK 74078-1058 msepans@math.okstate.edu

Slav Sharapov Department of Mathematics, North Dakota State Univ, Fargo, ND 58105

R. Showalter Department of Mathematics C1200, Univ of Texas at Austin, Austin, TX 78712 show@math.utexas.edu

Cesar Silva Department of Mathematics, Williams College, Williamstown, MA 01267 Cesar.E.Silva@williams.edu

Alexander Soshnikov Department of Mathematics, Princeton University, Fine Hall, Washington Road, Princeton, NJ 08544 soshnikv@math.princeton.edu

Birgit Speh Department of Mathematics, Cornell University speh@math.cornell.edu Mark Steinberger Department of Mathematics, University at Albany, Albany, NY 12222 mark@csc.albany.edu

Walter Strauss Brown University, Providence, RI

Catherine Sulem Department of Mathematics, University of Toronto, Toronto, Ontario, Canada

Thomas Ward Department of Mathematics, University of East Anglia, Norwich, NR6 5LB, England t.ward@uea.ac.uk

Reem Yassawi Department of Mathematics, McGill University, Montreal, Quebec, Canada reem@scylla.math.mcgill.ca

Andre Zelevinski Department of Mathematics, Northeastern University